# Classical optimization with imaginary time block encoding on quantum computers: The MaxCut problem

### v1

@@ Abstract
Finding ground state solutions of diagonal Hamiltonians is relevant for both theoretical as well as practical problems of interest in many domains such as finance, physics and computer science. These problems are typically very hard to tackle by classical computing and quantum computing could help in speeding up computations and efficiently tackling larger problems. Here we use imaginary time evolution through a new block encoding scheme to obtain the ground state of such problems and apply our method to MaxCut as an illustration. Our method, which for simplicity we call ITE-BE, requires no variational parameter optimization as all the parameters in the procedure are expressed as analytical functions of the couplings of the Hamiltonian. We demonstrate that our method can be successfully combined with other quantum algorithms such as quantum approximate optimization algorithm (QAOA). We find that the QAOA ansatz increases the post-selection success of ITE-BE, and shallow QAOA circuits, when boosted with ITE-BE, achieve better performance than deeper QAOA circuits. For the special case of the transverse initial state, we adapt our block encoding scheme to allow for a deterministic application of the first layer of the circuit.
@@ I Introduction
The advent of quantum computers has spurred a wide range of research into quantum advantage: problems that can be solved much faster, and/or at a much larger scale through quantum computing. Some of the most well known cases are integer prime factorization [1], and unstructured search [2], to name a few. An interesting venue of research is discrete optimization problems. These problems can be expressed as the minimization of a polynomial function of binary variables, and the cost function can be written as a diagonal Hamiltonian which encodes the optimal solution in the ground state [3, 4]. For instance, Quadratic Unconstrained Binary Optimization (QUBO) models several problems such as MaxCut, minimum vertex cover, and graph coloring [5, 6]. These problems have found applications in industry [7, 8], condensed matter physics [9], and beyond.
@@
Apart from special cases, finding the exact ground state of such diagonal Hamiltonians is usually difficult, and generally NP-hard [4, 3]. Often heuristic and approximate methods, such as simulated annealing [10] or semidefinite programming [11], are used to approximate solutions classically.
@@
Quantum computers could potentially find solutions beyond the realm of classical computation [12] as quantum hardware moves beyond the noisy intermediate-scale quantum (NISQ) era to fault tolerant computing [13]. Many quantum algorithms such as quantum approximate optimization algorithm (QAOA) [14, 15, 16, 17, 18], adiabatic evolution [19], or imaginary time evolution methods [20, 21, 22, 23] have been proposed to help determine the ground state. All these methods have their own drawbacks too, such as barren plateaus [24, 25] and other optimization related issues for QAOA [26], longer depth circuits for implementing adiabatic evolution[27] and larger number of measurements for successful post-selection of block encoded implementation of imaginary time evolution[28].
@@
QAOA has been widely explored recently in large sparse graph, and spin glasses [29], in particular in Sherrington-Kirkpatrick (SK) model of spin glass [30], on Maxcut problems [31] in both software and hardware [32, 33, 34]. In addition to two-body interactions, QAOA has also been applied to four-body interactions, such as the low autocorrelation binary sequences (LABS) problem [8].
@@
Imaginary time evolution, when applied to an initial state with non-zero overlap with the ground state, exponentially improves the overlap towards the ground state. Since it is a non-unitary operation, implementing it on a quantum computer is not trivial, and there has been several approaches in the literature. One approach is to find equivalent unitary evolutions, with appropriate quantum measurements and classical computations, such as in variational imaginary time evolution [20, 35] where a fixed ansatz is updated to follow McLachlan’s variational principle. Another approach discretizes the imaginary time evolution and approximates each step by a unitary constructed through tomography [21, 36, 37]. These methods require classical processing and multiple measurements during the time evolution which scale with both system and length of time. Other methods such as block-encoding schemes [38, 39, 40], encode the desired non-unitary operation in a bigger unitary matrix, which can then be implemented with ancilla qubits and post-selection. Recently an exact block-encoding scheme for discretized imaginary time evolution, based on unitary Restricted Boltzmann Machines (RBM), has been proposed [28] (for similar recent implementations see [23]).
@@
Here we utilize this imaginary time evolution block encoding (ITE-BE) scheme of the first order Trotter decomposition, to implement imaginary time evolution in solving classical optimization problems such as MaxCut. We also implement a new parametrization that allows to correct for failed post-selection when the initial state is in the traverse state. We find that ITE-BE is beneficial in providing solutions to MaxCut, and also as boosting technique to improve other methods such as QAOA. In Sec. II we define the MaxCut problem, QAOA and the block encoding method ITE-BE. Then, we proceed to perform numerical evaluations of these methods in Sec. III. In Sec. IV we summarize our findings and discuss their implications.
@@
Figure 1: Block-encoding-based circuit implementations for the imaginary-time propagator $e^{-\tau w_{jk}Z_{j}Z_{k}/2}$ with (a) a generic input state $|\psi\rangle$ and (b) the input state $|+\rangle_{j}\otimes|+\rangle_{k}$. Circuit parameters are given in eq. 8 and eq. 12. These block encodings can be applied for a generic Pauli string propagator $e^{-K\prod_{i}\sigma_{i}}$.
@@ II.1 The MaxCut Problem Definition
In this work, we focus on the MaxCut problem for a given undirected graph $G=(V,\tilde{E})$, where $V=\{1,\dotsc,N\}$ represents the set of vertices, and $\tilde{E}=\{(\langle j,k\rangle,w_{jk})\}$ denotes the set of edges with non-negative weights $w_{jk}\geq 0$ associated with the edge connecting vertices $j$ and $k$. The objective of this problem is to partition vertices of graph $G$ into two disjoint subsets such that the sum of weights of the edges linking the two subsets is maximized.
@@
However, it is NP-hard to design a classical algorithm that achieves $r\geq 16/17$ for MaxCut on all graphs [42]. This limitation motivates the exploration of quantum computing-based methods for solving the MaxCut problem, which are designed to maximize the quantum cost function $-\langle H_{C}\rangle$. The diagonal Hamiltonian $H_{C}$ for the MaxCut problem on graph $G$ is given by
@@
where $I$ is identity operator for $N$ qubit Hilbert space, $Z_{j}$ and $Z_{k}$ are Pauli-Z operators acting on qubits corresponding to vertices $j$ and $k$. Ideally, we expect the quantum algorithm to return a quantum state $|\psi\rangle$ as a superposition of computational basis $|\vec{z}\rangle$, where $\vec{z}\in\{0,1\}^{N}$ is a binary string representing a solution of the problem. In this context, each state $|\vec{z}\rangle$ is an eigenvector of $Z_{j}Z_{k}$ with eigenvalue ±$\pm 1$ for any edge $\langle j,k\rangle$, and the quantum cost function $C(\vec{z})=-\langle\vec{z}|H_{C}|\vec{z}\rangle$ has a similar structure to the classical cost function $C(\vec{s})$ in the integer quadratic programming problem.
@@ II.2 Quantum Approximate Optimization Algorithm
To find the optimal parameters $(\vec{\beta},\vec{\gamma})$, one typically starts from an initial guess and then performs numerical optimization to maximize $\langle H_{C}(\vec{\beta},\vec{\gamma})\rangle$ using feedback from the quantum computer. This iterative process continues until the maximum cost function value is found. A good choice of initial parameters can significantly reduce the computational cost of optimization. Heuristic strategies have been proposed to optimize $\langle H_{C}(\vec{\beta},\vec{\gamma})\rangle$ with respect to parameters $(\vec{\beta},\vec{\gamma})$ using a good initial guess [33].
Recently, Ref. [30, 29] applied QAOA to the Sherrington-Kirkpatrick (SK) model and derived near-optimal parameters $(\vec{\beta},\vec{\gamma})$ for $p<20$, and at $p=11$ they find that QAOA outperforms the standard semi-definite programming algorithm.
@@ II.3 ITE-BE Method
|$\ket{\psi(\tau)}=e^{-\tau H}\ket{\psi(0)},\tau>0.$ (5)
As long as the initial state |$\ket{\psi(0)}$ is not orthogonal to the ground state and $\tau$ is sufficiently large.
@@
In this approach, data will be discarded unless the ancilla qubit is measured to be $|0\rangle$. If the input state to a certain circuit block is $|+\rangle_{j}\otimes|+\rangle_{k}$, we can modify the block encoding parameters to deterministically apply the imaginary time propagator, see Fig. 1(b). We refer readers to Appendix A for further details.
@@ III Protocols and Evaluation
In this section, we design two protocols built upon the imaginary-time evolution to solve the MaxCut problem. The first protocol is a purely ITE-BE method starting with initial state $|+\rangle^{\otimes N}$, in which we adopt the circuit shown in Fig. 1(b) for the first layer and apply the circuit in Fig. 1(a) for the subsequent layers. This choice improves the post-selection rate by effectively setting to $1$ for the first layer, even when the ancilla is not measured to be|$\ket{0}$ state. The second protocol starts with the QAOA circuit, followed by ITE-BE, see Fig. 2. In essence, this method uses a QAOA circuit to prepare an initial state with a non-minimal overlap with the ground state. Then, the state is evolved through ITE-BE to significantly enhance the overlap.
@@
Both methods can be applied to general settings. As proof of concept, we numerically evaluate the performance of these protocols on unweighted 3-regular (u3R) graphs with vertex number $6\leq N\leq 12$. We finally report the $99.7\%$ uncertainty intervals for approximation ratio $r$, the probability of obtaining the optimal solution $p_{\rm opt}=N_{\rm opt}/N_{\rm tot}$ and the post-selection rate. Here, $N_{\rm tot}$ is the number of measurements made on data qubits after the post-selection of the ITE-BE method, and $N_{\rm opt}$ is the number of times one of the optimal solutions is obtained. Both metrics estimate the average quality of an algorithm, but $p_{\rm opt}$ is more conservative because it directly reflects the algorithm’s ability to find exact solutions. The post-selection rate denotes the success rate of the block encoding procedure, which is the probability in getting all |$\ket{0}$ state for the ancilla qubits.
@@ III.1 ITE-BE Only
Figure 3: The averaged approximation ratio $r$, probability of obtaining optimal solution $p_{\rm opt}$, and the post-selection success rate for randomly generated unweighed $3-$regular graphs within three standard deviations of the mean for ITE-BE method. This method guarantees convergence to the optimal solution, while the post-selection success rate decays exponentially. The convergence becomes slower and the success rate decreases when the problem size $N$ increases.
@@
Our first method is purely based on the imaginary time evolution. Since all $Z_{j}Z_{k}$ terms commute in the sum on the left-hand side of Eq. 7, we can rearrange the order of multiplication on the right-hand side of the same equation and separate them into two parts $V_{1}$ and $V_{2}$. For the first part, we define $V_{1}=\prod_{\langle j,k\rangle}e^{-\tau\frac{w_{jk}}{2}Z_{j}Z_{k}}$ using edges $\langle j,k\rangle$ that form a maximal matching in the graph, i.e., a set of disjoint edges without common vertices that includes as many nodes as possible. All imaginary propagators in $V_{1}$ can be implemented by the circuit with correction shown in Fig. 1(b) when the protocol starts with initial state $|+\rangle^{\otimes N}$. The second part consists of all remaining edges in the graph, and they can be only implemented using the circuit in Fig. 1(a) as the input state of each block is no longer $|+\rangle_{j}\otimes|+\rangle_{k}$. We have verified that this choice enhances post-selection success. To reduce the overhead of this approach, ancilla qubits for each block can be reused by performing mid-circuit measurements and resetting them to $|0\rangle$. In this case, the current shot terminates and a new shot begins if any of the ancilla qubits in the second part is measured to be $|1\rangle$. All data qubits are measured only after all circuit blocks for the imaginary time propagator $e^{-\tau H_{C}}$ have been executed.
@@
We then apply this method to solve the MaxCut problem on 100 randomly generated u3R graph instances with fixed $N$, where $w_{jk}=1$ for any $j,k$ in edges $E$ and each node is connected exactly to three other nodes. The imaginary time $\tau$ is varied from 00 to $2$ which allows us to observe how the results evolve until they are sufficiently close to the ideal solutions. For each $\tau$, we run the ITE-BE circuit with $10^{5}$ shots for each individual graph and calculate the approximation ratio, the probability of obtaining the optimal solution, and the post-selection success rate using data from successful executions. We repeat this process 10 times to accumulate enough samples and reduce statistical fluctuations. Finally, we exclude graphs with at least one overall failure (i.e., all $10^{5}$ shots discarded) in 10 repetitions, and report $99.7\%$ uncertainty intervals of above three metrics in Fig. 3. For comparison, we also show the bound of approximation ratio $r\approx 0.9326$ of classical semidefinite programming for u3R graphs [41]. All quantum circuits were simulated with the AerSimulator backend provided by the Qiskit Python package [43].
@@
Our simulation shows that, as expected, both the approximation ratio $r$ and the probability of obtaining optimal solutions $p_{\rm opt}$ increase monotonically with imaginary time $\tau$ across all graph sizes and eventually saturate near the optimal value of $1.0$ as $\tau$ approaches $2$. Meanwhile, the post-selection success rate decreases exponentially with $\tau$, which is consistent with the prediction of the success rate in Ref. [28]. We note that smaller graphs converge more quickly to the ground state and exhibit higher post-selection success rates compared to larger graphs due to fewer terms in the Hamiltonian. As the figure shows, $r\geq 0.5$ even at $\tau=0$ (just the $|+\rangle^{\otimes N}$ state). For further details on why this is the case, see Appendix D.

### v2

@@ Abstract
Optimization problems in finance, physics and computer science are typically very hard to tackle in classical computing and quantum computing could help speed up computations and provide efficient methods for tackling large problems. Typically, to treat the problem with a quantum computer, the optimal solution is cast as the ground state of a diagonal Hamiltonian. We develop a new method, called ITE-BE, based on a recent imaginary time algorithm, which requires no variational parameter optimization as all parameters can be derived analytically from the target Hamiltonian. We also demonstrate that our method can be successfully combined with other quantum algorithms such as quantum approximate optimization algorithm (QAOA). For illustration, here we study the MaxCut problem. We find that the QAOA ansatz increases the post-selection success of ITE-BE, and shallow QAOA circuits, when boosted with ITE-BE, achieve better performance than deeper QAOA circuits. For the special case of the transverse initial state, we adapt our block encoding scheme to allow for a deterministic application of the first layer of the circuit.
@@ I Introduction
The advent of quantum computers has spurred a wide range of research into quantum advantage: problems that can be solved much faster, and/or at a much larger scale through quantum computing. Some of the most well known cases are integer prime factorization [1], and unstructured search [2], to name a few. An interesting venue of research is discrete optimization problems [3]. These problems can be expressed as the minimization of a polynomial function of binary variables, and the cost function can be written as a diagonal Hamiltonian which encodes the optimal solution in the ground state [4, 5]. For instance, Quadratic Unconstrained Binary Optimization (QUBO) models several problems such as MaxCut, minimum vertex cover, and graph coloring [6, 7]. These problems have found applications in industry [8, 9], condensed matter physics [10], and beyond.
@@
Apart from special cases, finding the exact ground state of such diagonal Hamiltonians is usually difficult, and generally NP-hard [5, 4]. Often classical heuristic and approximate methods, such as simulated annealing [11] or semi-definite programming [12], are used to approximate solutions. Simulated annealing motivated by the annealing process in statistical mechanics, is a probabilistic method where an appropriate temperature and annealing schedule is defined to ‘cool’ the system, while the new configuration is accepted with an appropriate acceptance probability.
Semi-definite programming (SDP) minimizes a linear cost function subject to a semi-definite constraint, resulting in convex optimization problem which can be solved efficiently [13]. SDP could be applied by relaxing the original optimization problems (e.g., the MaxCut problem), followed by an appropriate rounding scheme to obtain approximate solutions [12].
@@
Quantum computers could potentially find solutions beyond the realm of classical computation [3] as quantum hardware moves beyond the noisy intermediate-scale quantum (NISQ) era to fault-tolerant computing [14]. Many quantum algorithms such as quantum approximate optimization algorithm (QAOA) [15, 16, 17, 18, 19], adiabatic evolution [20], or imaginary time evolution methods [21, 22, 23, 24] have been proposed to help determine the ground state. All these methods have their own drawbacks too, such as barren plateaus [25, 26] and other optimization related issues for QAOA [27], longer depth circuits for implementing adiabatic evolution[28] and larger number of measurements for successful post-selection of block encoded implementation of imaginary time evolution[29].
@@
QAOA has been widely explored recently in large sparse graphs, and spin glasses [30], in particular in Sherrington-Kirkpatrick (SK) model of spin glass [31], on MaxCut problems [32] in both software and hardware [33, 34, 35]. In addition to two-body interactions, QAOA has also been applied to four-body interactions, such as the low autocorrelation binary sequences (LABS) problem [9].
@@
Imaginary time evolution, otherwise known as Wick rotation [36], connects Euclidean and Minkowski space by performing an analytical continuation from real to imaginary time (see figure 6.1 from [37]), and has been used extensively in auxiliary field diffusion Monte Carlo studies [38, 39, 40]. This approach exponentially improves the overlap towards the ground state when applied to an initial state with non-zero overlap with the ground state. Since it is a non-unitary operation, implementing it on a quantum computer is not trivial, and there has been several approaches in the literature. One approach is to find equivalent unitary evolutions, with appropriate quantum measurements and classical computations, such as in variational imaginary time evolution [21, 41, 42] where a fixed ansatz is updated to follow McLachlan’s variational principle. Another approach discretizes the imaginary time evolution and approximates each step by a unitary constructed through tomography [22, 43, 44]. These methods require classical processing and multiple measurements during the time evolution which scale with both system and length of time. Other methods such as block-encoding schemes [45, 46, 47], encode the desired non-unitary operation in a bigger unitary matrix, which can then be implemented with ancilla qubits and post-selection. Recently an exact block-encoding scheme for discretized imaginary time evolution, based on unitary Restricted Boltzmann Machines (RBM), has been proposed [29] (for similar recent implementations see [24]).
@@
Here we utilize this imaginary time evolution block encoding (ITE-BE) scheme of the first order Trotter decomposition, to implement imaginary time evolution in solving classical optimization problems such as MaxCut. We also implement a new parametrization that allows to correct for failed post-selection when the initial state is in the traverse state. We find that ITE-BE is beneficial in providing solutions to MaxCut, and also as a boosting technique to improve other methods such as QAOA. In Sec. II we define the MaxCut problem, QAOA and the block encoding method ITE-BE. Then, we proceed to perform numerical evaluations of these methods in Sec. III. In Sec. IV we summarize our findings and discuss their implications.
@@
Figure 1: Block-encoding-based circuit implementations for the imaginary-time propagator $e^{-\tau w_{jk}Z_{j}Z_{k}/2}$ with (a) a generic input state $|\psi\rangle$ and (b) the input state $|+\rangle_{j}\otimes|+\rangle_{k}$. Circuit parameters are given in Eq. (8) and Eq. (12). These block encodings can be applied for a generic Pauli string propagator $e^{-K\prod_{i}\sigma_{i}}$.
@@ II.1 The MaxCut Problem Definition
[…]
@@
Quantum computing-based methods can also solve the MaxCut problem, which are designed to search for a quantum state that maximize the quantum cost function $\langle H_{C}\rangle$. The diagonal Hamiltonian $H_{C}$ for the MaxCut problem on graph $G$ is given by
@@
where $I$ is identity operator for $N$ qubit Hilbert space, $Z_{j}$ and $Z_{k}$ are Pauli-Z operators acting on qubits corresponding to vertices $j$ and $k$. Ideally, we expect the quantum algorithm to return a quantum state $|\psi\rangle$ as a superposition of computational basis $|\vec{z}\rangle$, where $\vec{z}\in\{0,1\}^{N}$ is a binary string representing a solution of the problem. In this context, each state $|\vec{z}\rangle$ is an eigenvector of $Z_{j}Z_{k}$ with eigenvalue $\pm 1$ for any edge $\langle j,k\rangle$, and the quantum cost function $C(\vec{z})=-\langle\vec{z}|H_{C}|\vec{z}\rangle$ has a similar structure to the classical cost function $C(\vec{s})$ in the integer quadratic programming problem.
@@ II.2 Quantum Approximate Optimization Algorithm
To find the optimal parameters $(\vec{\beta},\vec{\gamma})$, one typically starts from an initial guess and then performs numerical optimization to maximize $\langle H_{C}(\vec{\beta},\vec{\gamma})\rangle$ using feedback from the quantum computer. This iterative process continues until the maximum cost function value is found. A good choice of initial parameters can significantly reduce the computational cost of optimization. Heuristic strategies have been proposed to optimize $\langle H_{C}(\vec{\beta},\vec{\gamma})\rangle$ with respect to parameters $(\vec{\beta},\vec{\gamma})$ using a good initial guess [34]. Moreover, instead of searching on an instance-by-instance basis, the optimal parameters can be determined in advance for some special cases, including $p\leq 8$ QAOA for Sherrington-Kirkpatrick (SK) model [31] and $p\leq 17$ QAOA for MaxCut on large-girth $D$-regular graphs [49]. These values have been adopted in some software packages for QAOA simulation such as QOKit from JPMorgan Chase [50].
@@ II.3 ITE-BE Method
[…]
@@
In this approach, data will be discarded unless the ancilla qubit is measured to be $|0\rangle$. If the input state to a certain circuit block is $|+\rangle_{j}\otimes|+\rangle_{k}$, we can modify the block encoding parameters to deterministically apply the imaginary time propagator, see Fig. 1(b). In particular, the last $\rm CNOT$ gate could “fix”the data qubit if ancilla is measured to be $|1\rangle$, so no post-selection is needed for this circuit block. We refer readers to Appendix A for further details. For a given undirected graph $G=(V,E)$, the ITE-BE method requires $|V|+1$ qubits if reusing ancilla qubit for each block in Fig. 1 and a circuit with $O(|E|)$ depth according to Eq. (7).
@@ III Protocols and Evaluation
In this section, we design two protocols built upon the imaginary-time evolution to solve the MaxCut problem. The first protocol is a purely ITE-BE method starting with initial state $|+\rangle^{\otimes N}$. In this case, we adopt the circuit shown in Fig. 1(b) for the first layer, which improves the post-selection rate, and apply the circuit in Fig. 1(a) for the subsequent layers. The second protocol starts with the QAOA circuit, followed by ITE-BE, see Fig. 2. In essence, this method uses a QAOA circuit to prepare an initial state with a non-minimal overlap with the ground state. Then, the state is evolved through ITE-BE to significantly enhance the overlap.
@@
[…]
@@ III.1 ITE-BE Only
Figure 3: The averaged approximation ratio $r$, probability of obtaining optimal solution $p_{\rm opt}$, and the post-selection success rate for randomly generated unweighed 33-regular graphs within three standard deviations of the mean for ITE-BE method. Here, the approximation ratio starts from $0.5$ at $\tau=0$, which is expected according to Section 6.2.1 of Ref. [51]. This method guarantees convergence to the optimal solution, while the post-selection success rate decays exponentially. The convergence becomes slower and the success rate decreases when the problem size NN increases.
@@
Our first method is purely based on the imaginary time evolution. Since all $Z_{j}Z_{k}$ terms commute in the sum on the left-hand side of Eq. (7), we can rearrange the order of multiplication on the right-hand side of the same equation and separate them into two parts $V_{1}$ and $V_{2}$. For the first part, we define $V_{1}=\prod_{\langle j,k\rangle}e^{-\tau\frac{w_{jk}}{2}Z_{j}Z_{k}}$ using edges $\langle j,k\rangle$ that form a maximal matching in the graph, i.e., a set of disjoint edges without common vertices that includes as many nodes as possible. All imaginary propagators in $V_{1}$ can be implemented by the circuit with correction shown in Fig. 1(b) when the protocol starts with initial state $|+\rangle^{\otimes N}$. The second part consists of all remaining edges in the graph, and they can be only implemented using the circuit in Fig. 1(a) as the input state of each block is no longer $|+\rangle_{j}\otimes|+\rangle_{k}$. We have verified that this choice enhances post-selection success. To reduce the overhead of this approach, ancilla qubits for each block can be reused by performing mid-circuit measurements and resetting them to $|0\rangle$. In this case, the current shot terminates and a new shot begins if any of the ancilla qubits in the second part is measured to be $|1\rangle$. All data qubits are measured only after all circuit blocks for the imaginary time propagator $e^{-\tau H_{C}}$ have been executed.
@@
We then apply this method to solve the MaxCut problem on 100 randomly u3R graph instances generated by networkx Python package with fixed $N$, where $w_{jk}=1$ for any $j,k$ in edges $E$ and each node is connected exactly to three other nodes. The imaginary time $\tau$ is varied from 0 to $2$ which allows us to observe how the results evolve until they are sufficiently close to the ideal solutions. For each $\tau$, we run the ITE-BE circuit with $10^{5}$ shots for each individual graph and calculate the approximation ratio, the probability of obtaining the optimal solution, and the post-selection success rate using data from successful executions. We repeat this process 10 times to accumulate enough samples and reduce statistical fluctuations. Finally, we exclude graphs with at least one overall failure (i.e., all $10^{5}$ shots discarded) in 10 repetitions, and report $99.7\%$ uncertainty intervals of above three metrics in Fig. 3. All quantum circuits were simulated with the AerSimulator backend provided by the Qiskit Python package [52].
@@
Our simulation shows that, as expected, both the approximation ratio $r$ and the probability of obtaining optimal solutions $p_{\rm opt}$ increase monotonically with imaginary time $\tau$ across all graph sizes and eventually saturate near the optimal value of $1.0$ as $\tau$ approaches $2$. Meanwhile, the post-selection success rate decreases exponentially with $\tau$, which is consistent with the prediction in Ref. [29]. Such behavior is typical of repeated applications of block encoding due to the post selection probability being less than 1. If one chooses to employ a separate auxiliary qubit per Pauli string in the Hamiltonian, and post select at the end of all propagator applications, then one can resort to amplitude amplification [53] to improve the post-selection rate. In addition to the increased cost in terms of the number auxiliary qubits, this additional step would also increase the overall circuit depth. We note that smaller graphs converge more quickly to the ground state and exhibit higher post-selection success rates compared to larger graphs due to fewer terms in the Hamiltonian.